OpenStudy (anonymous):

Choose the correct answer. Which expression, when evaluated, yields a true statement? A. 16 – 2m ≤ 5 + 2m when m = 3 B. C. 4x – 1 > 2x + 7 when x = 4 D.

OpenStudy (anonymous):

ok so the other ones cold not be copyed but i was thinking C

OpenStudy (anonymous):

What happened to answer choices 'B' and 'D'?

OpenStudy (anonymous):

just told you

OpenStudy (anonymous):

Yeah, I saw it too late.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Neither A nor C is right.

OpenStudy (anonymous):

Substitute 4 for x.

OpenStudy (anonymous):

The statements have to be true and those aren't. What happened to B and D? I didn't understand what you said.

OpenStudy (anonymous):

4x-x/3greater than or = to 13 when x=3

OpenStudy (anonymous):

would this be right?

OpenStudy (anonymous):

4(4) -1 > 2(4) + 7 16 -1 > 8 + 7 15 > 15 Obviously 15 is not greater than 15, C is false.

OpenStudy (anonymous):

That one isn't right either.

OpenStudy (anonymous):

11 isn't greater than or equal to 13.

OpenStudy (anonymous):

dang ok thanks thats all oyu need to do dont stress it. :DD

OpenStudy (anonymous):

Lol okay. :)

OpenStudy (anonymous):

(4(3) - 3) / 3 >= 13 (12 - 3) /3 >= 13 (9 / 3) >= 13 3 >= 13 False.

OpenStudy (anonymous):

Option A: 16 – 2(3) ≤ 5 + 2(3) 16 - 6 <= 5 + 6 10 <= 11 A is true.

OpenStudy (anonymous):

@speartonion The answer is A.

OpenStudy (anonymous):

ohh ok i will go back then and correct it

OpenStudy (anonymous):

I don't think A is right. 11 is not equal to 10.

OpenStudy (anonymous):

@200205650 it's less than or equal to, meaning that it can be = to 11 or below

OpenStudy (anonymous):

Oh okay then I thought it had to be both. Sorry about that! :/ @ab854

OpenStudy (anonymous):

lol